OpenStudy (anonymous):

Writing and solving systems of equations: Write two equations for each problem, then solve. The sum of 6 times Petra's age and 8 times Kathy's age is 162. Kathy is 1 year more than twice as old as Petra is. Find each of their ages.

OpenStudy (anonymous):

6x(petra's)+8(kathy's)=163 ----->(1) kathy's=1+2(petra's) ----->(2)

OpenStudy (anonymous):

@chococatpiano

OpenStudy (anonymous):

Can you show the work for me?

OpenStudy (skullpatrol):

6p+8k=162 k=1+2p 6p+8(1+2p)=162 6p+8+16p=162 22p=154 p=7 k=1+2(7)=15 Petra is 7 years old Kathy is 15 years old @chococatpiano